How much heat must be added to a 0.3^m3 rigid volume containing water at 200°C in order that the final temperature be raised to 800°C? The initial pressure is 1 MP.
Answer #1

The mass of water contained in the tank is calculated:

m = P * V * MM / R * T = 9.87 atm * 300 L * 18 g / mol / 0.082 atm * L / mol * K * 473 K = 1374.16 g

The heat necessary to heat said mass is calculated:

q = m * Cp * ΔT = 1374.16 g * 0.00418 kJ / mol * ° C * (800 - 200) ° C = 3446. 4 kJ

If you liked the answer, please rate it in a positive way, you would help me a lot, thank you.